OpenStudy (anonymous):

6x-2x+11=-5 what is x????? please help

OpenStudy (kamille):

6x-2x+11=-5 6x-2x=-5-11 4x=-16|:4 x=-4

OpenStudy (anonymous):

Check: 6(-4)-2(-4)+11=-5 -24+8+11=-5 -5=-5 True